An interior decorator is buying wall paper for two square walls.One wall is 3ft. wider than the other.If she needs 89sq.ft of wall paper.What are the dimension of the wall?
Answer by jorel1380(3719)

Let w and w+3 be the one side of the 1st and 2nd walls, respectively. Then:
w^2 + (w+3)^2=89
2w^2+6w-80=0
w^2+3w-40=0
Solve for w and w+3.
🙂🙂🙂🙂